the perimeter of a triangle measures 62cm. the 2nd side is 2 times longer than the fist, and the 3rd side is 12cm larger than the second. What is the length of the longest side.
Answer by Paul(988)

Let the first side be x
Let the second side be y
Let the third side be z
x+y+z=62 (equation 1) (1)
y = 2x---->2x-y=0 (equation 2) (2)
z = 12+2x ----> 2x-z=-12 (equation 3) (3)
Eliminate z
x+y+z=62 (1)
2x-z=-12 (3)
-------- --> add the following
3x+y=50 (set 1)
Add set 1 to equation 2
3x+y=50
2x-y=0
5x=50
x=10

y=2(10)
y=20

z=2(10)+12
z=32

Hence, the first side 10 cm, the second side is 20 cm, and the third side is 32 cm.
Paul.